Both Bond Alpha and Bond Beta have 8% coupons, $1,000 face-value, make semi-annual payments and are priced at par value. Bond Alpha has 2 years to maturity, whereas Bond Beta has 15 years to maturity.

  1. If interest rates suddenly rise by 2%, what is the percentage change in the price of Bond Alpha relative to the original price?
  2. If rates were to suddenly fall by 2%, what would the percentage change in the price of Bond Alpha relative to the original price?
  3. If interest rates suddenly rise by 2%, what is the percentage change in the price of Bond Beta relative to the original price?
  4. If rates were to suddenly fall by 2%, what would the percentage change in the price of Bond Beta relative to the original price?
Answer #1

a

Part 1
Change in YTM =2
Bond Alpha
                  K = Nx2
Bond Price =∑ [( Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =2x2
Bond Price =∑ [(8*1000/200)/(1 + 10/200)^k]     +   1000/(1 + 10/200)^2x2
                   k=1
Bond Price = 964.54
%age change in price =(New price-Old price)*100/old price
%age change in price = (964.54-1000)*100/1000
= -3.55%
b
Bond Beta
                  K = Nx2
Bond Price =∑ [( Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =15x2
Bond Price =∑ [(8*1000/200)/(1 + 10/200)^k]     +   1000/(1 + 10/200)^15x2
                   k=1
Bond Price = 846.28
%age change in price =(New price-Old price)*100/old price
%age change in price = (846.28-1000)*100/1000
= -15.37%
c
Part 2
Change in YTM =-2
Bond Alpha
                  K = Nx2
Bond Price =∑ [( Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =2x2
Bond Price =∑ [(8*1000/200)/(1 + 6/200)^k]     +   1000/(1 + 6/200)^2x2
                   k=1
Bond Price = 1037.17
%age change in price =(New price-Old price)*100/old price
%age change in price = (1037.17-1000)*100/1000
= 3.72%
d
Bond Beta
                  K = Nx2
Bond Price =∑ [( Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =15x2
Bond Price =∑ [(8*1000/200)/(1 + 6/200)^k]     +   1000/(1 + 6/200)^15x2
                   k=1
Bond Price = 1196
%age change in price =(New price-Old price)*100/old price
%age change in price = (1196-1000)*100/1000
= 19.6%